# Deep Probabilistic Direction Prediction in 3D with Applications to Directional Dark Matter Detectors

### v1

@@ 4.4 Performance
To verify that the model is robust against modest variation of the simulation parameters, we create a new test set that is equivalent to the 40 keV test set with 443 um diffusion, but simulated at $22^{\circ}$ instead of $20^{\circ}$. Even though the HCN model is trained on $20^{\circ}$ simulations, the cosine distance loss it achieves on the $20^{\circ}$ and $22^{\circ}$ test sets agrees within 0.48%. This indicates that small fluctuations in temperature have a very small systemic effect on our model.
@@ 5 Conclusion
This work integrates all the essential components required for applying deep learning to modern directional recoil detectors. For the first time, we describe a deep probabilistic approach for predicting 3D direction. Our model utilizes the von Mises-Fisher to predict distributions on $\mathbb{S}^{2}$ and we specify approximations of the NLL that enable stable training. Additionally, we develop and perform two tests to assess model calibration and we demonstrate that rejecting events with high estimated uncertainty improves accuracy. We also develop a sparse 3D convolution approach for processing high-dimensional event data in Gas TPC detectors. Together, these contributions provide a probabilistic deep learning approach that advances the state-of-the-art for analyzing directional recoil detector data. Moreover, the method is applicable to any task requiring 3D direction estimation with uncertainty.
The benefits of obtaining directional information are profound and could prove game-changing in several areas of fundamental physics. Most importantly, dark matter detection experiments could reject solar neutrino backgrounds based on recoils pointing back towards the sun and positively identify dark matter from the observed angular distribution of recoils (Billard et al., 2014; O’Hare, 2021). This would overcome two major limitations in current experiments seeking to identify the particle nature of dark matter, which is arguably the most urgent problem in contemporary physics. Directional measurements of electron recoils induced by solar neutrinos would enable improved measurements of neutrinos from the Sun’s “CNO cycle” which could settle a long-standing puzzle known as the solar abundance problem (Villante and Serenelli, 2019; O’Hare et al., 2022). Directional measurements of photoelectrons produced by X-rays enable us to map the X-ray polarization of extended astrophysical sources, as in the Imaging X-ray Polarimetry Explorer  (Weisskopf et al., 2016). A comprehensive review of directional recoil detection and its benefits can be found in  Vahsen et al. (2021).
@@ Appendix B Application to a Toy Problem
In this section, we provide an illustrative example of our method on a toy problem of determining the direction of 3D arrows. Arrows with a random direction, fixed shape, and uniform density are generated and binned into a $(120,120,120)$ voxel grid. An example is displayed in figure 8. We generate $6000$ arrows, $4000$ are used for the training set and $1000$ are used for the validation and test sets. The RCN and HCN models are trained on the training set while their validation loss is used for early stopping. If the validation loss has not decreased in the last 2 epochs, training is stopped. The weights corresponding to the lowest validation loss are saved and used for the final model.
In this simple case, both models are fully trained within minutes. The cosine distance loss of the RCN and HCN models on the test data set is $4\times 10^{-4}$ and $8\times 10^{-4}$, respectively. As expected, the direction predictions are much more accurate than Section 4. The lowest predicted $\kappa_{i}$ by HCN on the test data set is $443$, meaning the model is much more confident in the predicted directions than for the electron recoil case. The code needed to create the arrow data sets, train our models, and test our models is available at https://github.com/mghrear/3D_Heteroscedastic_Convnet.

### v2

@@ 4.4 Performance
Since our model is trained over a large range of diffusion and the simulated electrons include a significant amount of variability in terms of track shape and the amount of deposited ionization, we expect our model to be robust against modest variations in simulation parameters, including environmental conditions. We verify this with a test on varying temperature, which is the least tightly controlled environmental parameter in most setups. We create a new test set that is equivalent to the 40 keV test set with 443 um diffusion, but simulated at $22^{\circ}$ instead of $20^{\circ}$. Even though the vMF-NN model is trained on $20^{\circ}$ simulations, the cosine distance loss it achieves on the $20^{\circ}$ and $22^{\circ}$ test sets agrees within 0.48%. This indicates that small fluctuations in temperature have a very small systemic effect on our model.
4.5 Comparison to Gauss-NN
To highlight the advantages of the proposed vMF-NN model we compare to a second deep probabilistic model with the same architecture except it outputs the parameters of an isotropic 3D Gaussian distribution. This Gauss-NN demonstrates the effect of using the wrong manifold for direction detection. Instead of a distribution over the sphere, it outputs a Gaussian p.d.f. over 3D space. The target direction (a point on the unit sphere) is modeled as a point in 3D space sampled from this distribution. The Gaussian is constrained to be isotropic in order to facilitate comparison to the vMF, so a single scalar parameter $\sigma$ controls the variance of the predicted distribution centered on $\hat{\mathbf{y}}_{i}$ (which is constrained to be on the sphere). Training minimizes the NLL,
$NLL_{\textrm{Gauss}}=\sum_{i}^{N}log(\sigma_{i}^{3})+\frac{(\mathbf{y}_{i}-\hat{\mathbf{y}}_{i})^{2}}{2\sigma_{i}^{2}}.$ (7)
While this model is not a distribution over the sphere, it induces a distribution on the sphere known as the projected normal distribution in directional statistics (Wang and Gelfand, 2013). Optimizing the NLL of the induced projected normal distribution is not practical because computing the p.d.f. requires integration along the radial direction, but the Gauss-NN objective can be viewed as approximating that of the induced projected normal distribution on spherical data. Sampling from the induced projected normal distribution is trivial, and like the vMF-NN, cuts can be made on the $\sigma$ parameter to reject low-certainty samples. Thus, the Gauss-NN is an alternative to the vMF-NN with two key disadvantages: (1) computing the likelihood function over directions is computationally expensive, and (2) the approximate objective function contributes to miscalibrated uncertainty estimates.
To rigorously compare the performance of the deep learning models, we train each model five times with different random initializations and random mini-batches. The mean and standard deviation of the cosine distance loss on the two test sets are presented in Table 1. The two probabilistic models, vMF-NN and Gauss-NN, have nearly identical performance. The Det-NN has slightly lower error, but the difference is significant. The Best-Expected algorithm achieves the lowest error (but as we show above the vMF-NN can surpass it in performance with a 1% efficiency cut).
Model Name 40 keV, $443\,{\upmu}{\rm m}$ 50 keV, $200\,{\upmu}{\rm m}$
vMF-NN $0.10562\pm 0.00116$ $0.06362\pm 0.00064$
Gauss-NN $0.10540\pm 0.00095$ $0.06372\pm 0.00088$
Det-NN $0.10278\pm 0.00104$ $0.06236\pm 0.00042$
Best-Expected $0.098$ $0.0569$
Table 1: The mean and standard deviation of the cosine distance loss achieved by an ensemble of 5, for each of the models, on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing. Since Best-Expected does not have any weights that need to be initialized, it performance is captured by a single value.
The fact that the homoscedastic model performs slightly better than the heteroscedastic models is not surprising. This phenomena has been studied in the machine learning literature (Skafte et al., 2019; Seitzer et al., 2022) and there exist methods to reduce this gap (Stirn et al., 2023; Immer et al., 2023). The two probabilistic models performing similarly is explained by the models having identical architectures and very similar objective functions (Equation 4 and Equation 7). The main differentiator between the vMF-NN and Gauss-NN is the calibration — while the vMF-NN is shown to be well calibrated in Figure 6(a), the Gauss-NN suffers from systematic miscalibration in which the variance is overestimated (Figure 8). This is explained by the mismatch between the topology of the probabilistic model (in 3D) and the data (on the sphere). Together with the requirement of computing an integral to obtain the likelihood function for the induced projected normal distribution, this makes the Gauss-NN model a poor alternative compared to the vMF-NN for applications in which accurate calibration is needed. However, Figure 9 shows that the Gauss-NN performs just as well on our application, where only the relative ordering of the uncertainties is important for effectively rejecting low-certainty samples.
Figure 8: 2D histogram of predicted uncertainty in the Gauss-NN ($\sigma$) and the Euclidean distance ($r$) between predicted and target directions for all test data. The red curve is the mean $r$ for each bin in $\sigma$. The black dashed curve is the expectation when $r$ is sampled from a 3D Gaussian with standard deviation $\sigma$. The systematic bias demonstrates a miscalibration of the Gauss-NN model.
Figure 9: Cosine distance loss for the Gauss-NN and vMF-NN models at different efficiency cuts. Five models of each type were trained on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing. The mean and standard deviation of the test losses are plotted versus efficiency cuts where the $\sigma$ and $\kappa$ parameters are used to remove samples with high predicted uncertainty for the Gauss-NN and vMF-NN models, respectively.
5 Discussion
This work integrates all the essential components required for applying deep learning to modern directional recoil detectors. We describe a deep probabilistic approach for predicting 3D direction with well-calibrated uncertainties that can be used to reject low-certainty events. Using the design patterns of deep learning, we developed a 3D convolution approach that processes high-dimensional event data from Gas TPC detectors. Together, these contributions provide a probabilistic deep learning approach that advances the state-of-the-art for analyzing directional recoil detector data. Moreover, the method is applicable to any task requiring 3D direction estimation with uncertainty.
The benefits of obtaining directional information are profound and could prove game-changing in several areas of fundamental physics. Most importantly, dark matter detection experiments could reject solar neutrino backgrounds based on recoils pointing back towards the sun and positively identify dark matter from the observed angular distribution of recoils (Billard et al., 2014; O’Hare, 2021). This would overcome two major limitations in current experiments seeking to identify the particle nature of dark matter, which is arguably the most urgent problem in contemporary physics. Directional measurements of electron recoils induced by solar neutrinos would enable improved measurements of neutrinos from the Sun’s “CNO cycle” which could settle a long-standing puzzle known as the solar abundance problem (Villante and Serenelli, 2019; O’Hare et al., 2022; Lisotti et al., 2024). Directional measurements of photoelectrons produced by X-rays enable us to map the X-ray polarization of extended astrophysical sources, as in the Imaging X-ray Polarimetry Explorer  (Weisskopf et al., 2016). A comprehensive review of directional recoil detection and its benefits can be found in  Vahsen et al. (2021).
@@ 6 Conclusion
For the first time, we describe a deep probabilistic approach for predicting 3D direction. The model utilizes the von Mises-Fisher to predict distributions on $\mathbb{S}^{2}$ and we specify approximations of the NLL that enable stable training. Two tests are developed for assessing model calibration in direction prediction applications and the model is shown to be well-calibrated in our experiments.
The approach is demonstrated on an application to directional recoil detectors in physics. The model is compared to a variety of alternative approaches: a non-ML approach, a deterministic ML model without uncertainty, and a probabilistic deep learning model that does not account for the topology of the data. The results show that the ML models drastically out-perform the non-ML approach. While all the ML approaches perform very similarly in terms of cosine distance error, the probabilistic ML approach is shown to realize significantly lower errors by rejecting a small set of low-certainty events. Finally, it is argued that deep probabilistic direction prediction models that do not account for the true topology of the data have practical disadvantages and are less well-calibrated.
@@ Appendix B Application to a Toy Problem
In this section, we provide an illustrative example of our method on a toy problem of determining the direction of 3D arrows. Arrows with a random direction, fixed shape, and uniform density are generated and binned into a $(120,120,120)$ voxel grid. An example is displayed in figure 10. We generate $6000$ arrows, $4000$ are used for the training set and $1000$ are used for the validation and test sets. The Det-NN and vMF-NN models are trained on the training set while their validation loss is used for early stopping. If the validation loss has not decreased in the last 2 epochs, training is stopped. The weights corresponding to the lowest validation loss are saved and used for the final model.
In this simple case, both models are fully trained within minutes. The cosine distance loss of the Det-NN and vMF-NN models on the test data set is $4\times 10^{-4}$ and $8\times 10^{-4}$, respectively. As expected, the direction predictions are much more accurate than Section 4. The lowest predicted $\kappa_{i}$ by vMF-NN on the test data set is $443$, meaning the model is much more confident in the predicted directions than for the electron recoil case. The code needed to create the arrow data sets, train our models, and test our models is available at https://github.com/mghrear/3D_Heteroscedastic_Convnet.